OpenStudy (anonymous):

Calculate the # of moles present. 1.34kg of potassium bromide

OpenStudy (ookawaiioo):

1.34 kg --- 0.00134 grams/ MW of KBR = # of moles

sam (.sam.):

You can use the ratio method, For 1 mole of KBr you have 39+80 g of KBr \[1340 \text{g of KBr} \times \frac{1~\text{mole of KBr}}{39+80~\text{g of KBr}}\] \[1340 \cancel{\text{g of KBr}} \times \frac{1~\text{mole of KBr}}{39+80~\cancel{\text{g of KBr}}}=11.26 \text{ moles of KBr}\]

OpenStudy (ookawaiioo):

@.Sam. Well done :)